Output of good X Short run average costs (AVC) Average Revenue (AR)

1 110 300

2 95 250

3 80 210

4 75 180

5 82 150

6 85 120

7 90 100

8 100 90

9 110 80

10 120 70

Given the table above, calculate Total Variable costs (TVC), Total Costs (TC), Marginal costs (MC, Marginal revenue (MR) and profit (loss) at each level of output.determine the profit maximizing level of output. Calculate the smallest rise in TVC that would force the firm to cease production in the short run.distinguish between the short run and long run periods of production.





1
Expert's answer
2021-07-23T15:11:01-0400

TR"=" Price"\u00d7" Quantity

AR"=" P"(" Price")"

TVC"=" AVC"\u00d7" Q

TC"=" TVC"\u00d7" TFC


MC= "\frac{\Delta TC}{\Delta Q}" "=" "\frac{\Delta TVC}{\Delta Q}"


MR"=" "\frac{\Delta TR}{\Delta Q}"

Profit"\/" loss"=" TR"-" TC


Profit maximizing level output is the highest gap between TR and TC. Answer is 310


A firm should cease production when it incurs a loss. Therefore TVC above 510 will lead to a loss.


Short run production refers to a production cycle in which at least one factor is fixed. The factors include: labor, materials, equipment, capital and real property.

On the other hand long run production occurs when all factors of production fluctuate.